OpenStudy (anonymous):

Jay Downs only worked 15 of his scheduled hours last week. If he worked 75% of his scheduled hours, how many hours was he scheduled to work?

OpenStudy (anonymous):

15 : 75 x : 100 cross multiply

OpenStudy (anonymous):

15 is 75% (i.e. 3/4) of S, his scheduled hours 15=0.75S S=15/0.75=20

OpenStudy (amistre64):

75% is 3 out of 4 15/3 = 5 ; 5*4 = 20

OpenStudy (anonymous):

Thank you! Could you please show the process?

OpenStudy (anonymous):

everyone already did show the process. read it.

OpenStudy (anonymous):

? put 15 over 75 and x over 100? Im sorry I still dont understand. Im helping my sibling with basic math and forgot the process... Sorry!

OpenStudy (amistre64):

if 15 units is 75%; then divide 15 by 75 to determine 1%; the value at 1% is multiplied by 100 to get the total amount

OpenStudy (anonymous):

thank you!